The length of a rectangle is decreased by three feet, and the width is increased by one foot, forming a square region having an area of 25 square feet. What is the area of th original rectangular region?
Answer by Nate(3500)

If the square's area is 25 squared feet, then each side is 5 feet.
Length=5+3=8 feet
Width=5-1=4 feet